I am not sure the answer to this homework question. Please help with it.
10n=36+2n
-2n -2n
8n=36
8 8
n=4.5
or would it be n= no solution?

Answer by Earlsdon(6294)

You did everything right and your answer is correct.
You can always check your answer by substituting n = 4.5 into the given equation and see if the two sides are equal.
10n = 36+2n Substitute n = 4.5
10(4.5) = 36+2(4.5) Evaluate.
45 = 36+9
45 = 45 So your solution works!